OpenStudy (anonymous):

The value P of an initial investment of A dollars after 2 years is given by P = A(1 + r)2 , where r is the annual percentage rate earned by the investment. If an initial investment of $1300 grew to a value of $1544.53 in 2 years, what was the annual percentage rate?

OpenStudy (anonymous):

1544.53 = 1300 (1+r)^2 1544.53/1300 = (1+r)^2 1.1881 = (1+r)^2 1+ r = Sqrt[1.1881] r = Sqrt[1.1881]-1 = 0.09 9%